This is a multiple answer question   


let a and b be two non null vectors such that |a+b|=|a-2b| . Then the value of |a|/|b| May be : 


(a) 1/4  (b) 1/8  (c) 1 (d) 2.  Pls solve the question in detail

Grade:12th Pass

1 Answers

ruchi yadav
10 years ago
open the mod like this
a+b = +(a=2b)

i.e a+b = a-2b which is not possible because b is non null vector.
or
a+b = -(a-2b)

so a/b = 1/2


Thanks
Ruchi ,

Askiitians faculty